A student finishes the first half of a exam in 2/3 the time it takes him to finish the second half. If the entire exam takes him an hour, how many minutes does he spend on the first half of the exam?


I found out that 2/3 equals 40min. So I subtracted 60min-40 and got 20, but the book says its 24min. How do I get 24 as my answer?

Answer by checkley79(3341)

2/3X+X=60
2X/3+X=60
(2X+3X)/3=60
5X/3=60
5X=3*60
5X=180
X=180/5
X=36 TIME FOR THE SECOND HALF OF THE TEST.
2/3X=2/3*36=72/3=24 ANS. FOR THE FIRST HALF OF THE TEST.